OpenStudy (anonymous):

1,49,1625,3649,..... what is the next term in this series?

OpenStudy (anonymous):

6481

OpenStudy (anonymous):

ok

OpenStudy (accessdenied):

yeah, looks like you concatenate two consecutive squares together for each term 0^2 || 1^2 = 1 2^2 || 3^2 = 4 || 9 = "49" 4^2 || 5^2 = 16 || 25 = 1625